An object is released freely at high of 10 meters from the ground surface, what must be the velocity before it strikes the ground?


1
Expert's answer
2021-09-23T08:32:54-0400
"v^2=2gh\\v^2=2(9.8)(10)\\v=14\frac{m}{s}"